# SubZero: Random Subspace Zeroth-Order Optimization for Memory-Efficient LLM Fine-Tuning

### v1

@@ 6.1 Performance with Different Experimental Settings
6.2 Performance with Non-Differentiable Objectives
Following MeZO (Malladi et al., 2023), we respectively apply SubZero to fine-tune RoBERTa-large and OPT-13B using two non-differentiable objectives: accuracy and F1. As a baseline, we also report results using the cross-entropy objective with Adam. As shown in Table 4, SubZero consistently outperforms MeZO across both non-differentiable objectives and the cross-entropy benchmark, demonstrating its effectiveness across varying optimization goals.
@@ 6.3 Memory Usage and Wall-Clock Time Analysis
Table 5 compares the memory consumption and wall-clock time of ZO methods (MeZO and SubZero), SGD, and inference-only approaches (zero-shot and in-context learning (ICL)) using OPT-13B. Since inference-only methods do not involve fine-tuning, they have zero wall-clock time and their memory usage reflects only the inference load. For fine-tuning, all methods were run for 20K steps. The ZO methods, including SubZero, achieved over a 1.8× reduction in memory usage compared to SGD. Notably, SubZero’s memory footprint closely aligns with MeZO’s, while offering improved performance.
Although SubZero introduces additional computational overhead for generating projection matrices via QR decomposition, this extra time represents less than 5% of the total wall-clock time. It is important to note that due to differences in how steps are defined between ZO methods and SGD, direct wall-clock time comparisons between the two are not entirely meaningful.
Table 5: Memory usage (GB) and wall-clock time (minutes) of fine-tuning OPT-13B, with SGD’s batch size being 8 for SQuAD and 16 for other tasks.
Task SST-2 WIC SQuAD
Method Mem. Time Mem. Time Mem. Time
Zero-shot/ICL 24.2 0 24.8 0 27.2 0
SGD(FT) 48.9 190.3 48.9 257.3 122.7 623.7
MeZO(FT) 26.1 324.9 26.6 370.5 37.4 670.2
SubZero(FT) 26.5 337.3 27.1 385.3 37.8 690.5
MeZO(LoRA) 26.1 123.9 26.6 171.6 37.4 476.7
SubZero(LoRA) 26.1 130.3 26.6 179.7 37.4 486.5
@@ 6.4 Ablation Study
We conducted a thorough investigation of the effectiveness of our proposed techniques. Table 8 shows that using a column-orthogonal projection matrix significantly outperforms a Gaussian random projection matrix, primarily due to the low-rank structure of the perturbation matrices. This low-rank perturbation is key to improving the quality of gradient estimation.
Next, Table 8 explores the effects of subspace rank $r$ and update frequency $T_{0}$ in Algorithm 3. The results demonstrate that SubZero is robust to variations in the subspace rank. However, performance drops sharply when the update frequency is too low, as the optimization becomes constrained to a single subspace for too long, limiting its adaptability.
Table 6: Orthogonal or random projection matrix.
Dataset Ortho. Accuracy
RTE 67.5
74.0
WSC 59.6
65.1
Table 7: Subspace change frequency $T_{0}$ and rank $r$.
$T_{0}$$r$ 32 64 128
@@ 7 Conclusion
We have demonstrated that SubZero effectively fine-tunes large LLMs across various tasks and schemes with a memory cost comparable to that of inference. Extra experiments indicate that SubZero can optimize non-differentiable objectives. Our theory explains how SubZero reduces the variance of gradient estimates and accelerates convergence.
@@
Limitation. In addition to the SGD optimizer, we have yet to explore combining SubZero with other first-order optimizers, such as Adam. While SubZero is also compatible with other memory-efficient techniques like parameter quantization (Li et al., 2024), we have not thoroughly investigated the practical effects of these combinations. We will leave these explorations for future work.
Appendix A Appendix
A.1 Prompt Templates
For autoregressive LLMs, we have three task types: classification, multiple-choice, and question answering. We adopt the prompt templates for various tasks in (Malladi et al., 2023), which are summarized in Table 9. For masked LLMs, we also adopt the prompt templates in (Malladi et al., 2023) and present them in Table 10.
Table 9: The prompt templates used in the OPT-1.3B, OPT-13B, LLama2-7B, and Mistral-7B experiments.
Task Type Prompt
SST-2 cls. <text> It was terrible/great
RTE cls. <premise>
Does this mean that "<hypothesis>" is true? Yes or No?
Yes or No
CB cls. Does this mean that "<hypothesis>" is true? Yes or No?
Yes/No/Maybe
BoolQ cls. <passage> <question>?
Yes/No
WSC cls. <text>
In the previous sentence, does the pronoun "<span2>" refer to <span1>? Yes or No?
Yes/No
WIC cls. Does the word "<word>" have the same meaning in these two sentences? Yes, No?
<sentence1>
<sentence2>
Yes/No
MultiRC cls. <paragraph>
Question: <question>
I found this answer "<answer". Is that correct? Yes or No?
Yes/No
COPA mch. <premise> so/because <candidate>
ReCoRD mch. <passage>
<query>.replace("@placeholder", <candidate>)
SQuAD QA Title: <title>
Context: <context>
Question: <question>
Answer:
DROP QA Passage: <context>
Question: <question>
Answer:
Table 10: The prompt templates used in RoBERTa-large experiments. $C$ is the number of classification categories.
Task $C$ Type Prompt
SST-2 2 sentiment cls. <sentence1> It was great/terrible
SST-5 5 sentiment cls. <sentence1> It was great/good/okay/bad/terrible
MNLI 3 NLI <sentence1> ? Yes/Maybe/No , <sentence2>
SNLI 3 NLI <sentence1> ? Yes/Maybe/No , <sentence2>
A.2 Datasets
Following (Malladi et al., 2023), we use SuperGLUE (Wang et al., 2019) for OPT experiments, including BoolQ (Clark et al., 2019), CB (de Marneffe et al., 2019), COPA (Roemmele et al., 2011), MultiRC (Khashabi et al., 2018), ReCoRD (Zhang et al., 2018), RTE (Dagan et al., 2005; Bar Haim et al., 2006; Giampiccolo et al., 2007; Bentivogli et al., 2009), WiC (Pilehvar & Camacho-Collados, 2019), and WSC (Levesque et al., 2012). We also utilize SST-2 (Socher et al., 2013a) and two question answering (QA) datasets, SQuAD (Rajpurkar et al., 2016) and DROP (Dua et al., 2019).
For LLama2-7B and Mistral-7B, we use CB (de Marneffe et al., 2019) in the full-parameter tuning and three PEFT schemes. For OPT-1.3B, we utilize SST-2 (Socher et al., 2013a) in the full-parameter tuning and three PEFT schemes.
For RoBERTa-large, we consider classification datasets: SST-2 (Socher et al., 2013a), SST-5 (Socher et al., 2013a), MNLI (Williams et al., 2018), and SNLI (Bowman et al., 2015). Following Malladi et al. (2023), the test set has 1000 examples for fast iteration, while we have 512 examples per class for both training and validation.
A.3 Hyperparameters
Using a larger batch size can consistently reduce the variance in ZO optimization, thus enhancing fine-tuning performance (Malladi et al., 2023; Gautam et al., 2024; Yang et al., 2024). However, this increase in batch size also raises the time for forward passes and significantly elevates memory usage. We focus on developing ZO methods that minimize variance and improve performance with small batch sizes, with a default setting of 16. In some SGD experiments, like on MultiRC and SQuAD, the batch size is reduced to 8 due to limited GPU resources.
Consistent with previous studies (Malladi et al., 2023; Zhang et al., 2024; Liu et al., 2024; Yang et al., 2024), we employ SGD without momentum by default to maintain memory efficiency. SGD utilizes linear learning scheduling, while all ZO methods apply a constant learning rate, with weight decay set to 0.
For RoBERTa, we run Adam for 1K steps and ZO methods for 100K steps. In the rest experiments, we run Adam for 5 epochs and SGD and ZO methods for 20K steps.
We follow previous work to set the hyperparameters in the PEFT schemes (Malladi et al., 2023; Zhang et al., 2024). For LoRA, the rank is set to 8 and $\alpha$ is set to 16. For prefix tuning, the length of prefix tokens is set to 5, and we initialize these tunable representations by randomly sampling tokens from the vocabulary and then passing them through the LLM to get their keys and values at different attention layers. For prompt tuning, the length of prompt virtual tokens is set to 10, and the prompt tokens are initialized with actual token values from the model’s embedding.
We present the hyperparameter search grids in Tables 11 and 12 to assist with result reproduction. For OPT-1.3B, we utilize the same hyperparameter settings as in Table 12. For Roberta-large, we use a learning rate of {1e-6, 5e-6} and $\varepsilon$=1e-3 for MeZO and SubZero, with a batch size of 64. The rank for SubZero is set to {8, 16, 24}, and subspace change frequency is adjusted to {1000, 2000}.
Table 11: The hyperparameter search grids for OPT-13B. For each task, we run 20K steps for ZO methods (MeZO, S-MeZO, and SubZero) and 5 epochs for SGD. We record the best model checkpoint based on the validation loss every 500 training steps.
Experiment Hyperparameters Values
MeZO(FT) batch size 16
learning rate {1e-7, 2e-7, 5e-7, 1e-6}
$\varepsilon$ 1e-3
MeZO(LoRA) batch size 16
learning rate {1.5e-5, 3e-5, 5e-5}
$\varepsilon$ 1e-3
S-MeZO(FT) batch size 16
learning rate {1e-6, 5e-6}
$\varepsilon$ 1e-3
sparse rate 0.75
S-MeZO(LoRA) batch size 16
learning rate {5e-5, 1e-4, 1e-3}
$\varepsilon$ 1e-3
Sparse rate 0.75
SubZero(FT) batch size 16
learning rate {1e-7, 2e-7, 5e-7, 1e-6}
$\varepsilon$ 1e-3
rank {32, 64, 128, 256 }
subspace change frequency {500, 1000, 2000}
SubZero(LoRA) batch size 16
learning rate {1.5e-5, 3e-5, 5e-5}
$\varepsilon$ 1e-3
rank {4, 8, 16}
subspace change frequency {500, 1000, 2000}
SGD(FT) batch size 16
Learning rate {1e-4, 1e-3, 5e-3}
Table 12: The hyperparameter search grids for LLama2-7B and Mistral-7B. For each task, we run 20K steps for ZO methods (MeZO and SubZero) and 5 epochs for SGD. We record the best model checkpoint based on the validation loss every 500 training steps.
Experiment Hyperparameters Values
MeZO(FT) batch size 16
learning rate {1e-7, 5e-7, 1e-6}
$\varepsilon$ 1e-3
MeZO(LoRA) batch size 16
learning rate {1e-6, 5e-6, 1e-5, 3e-5}
$\varepsilon$ 1e-3
MeZO(Prefix) batch size 16
learning rate {1e-3, 5e-3, 1e-2}
$\varepsilon$ 1e-1
MeZO(Prompt) batch size 16
learning rate {1e-3, 5e-3, 1e-2}
$\varepsilon$ 1e-1
SubZero(FT) batch size 16
learning rate {1e-7, 5e-7, 1e-6}
$\varepsilon$ 1e-3
rank {24, 48}
subspace change frequency 1000
SubZero(LoRA) batch size 16
learning rate {1e-6, 5e-6, 1e-5, 3e-5}
$\varepsilon$ 1e-3
rank {4, 8}
subspace change frequency 1000
SubZero(Prefix) batch size 16
learning rate {1e-3, 5e-3, 1e-2}
$\varepsilon$ 1e-1
rank {4, 8}
subspace change frequency 1000
SubZero(Prompt) batch size 16
learning rate {1e-3, 5e-3, 1e-2}
$\varepsilon$ 1e-1
rank {16, 24}
subspace change frequency 1000
SGD(FT) batch size 16
Learning rate {1e-5, 1e-4, 1e-3, 5e-3}
A.4 Implementation Details
We use one A800 GPU with the PyTorch 2.1.0+CUDA 11.8 framework for ZO methods and, if needed, two A800 GPUs for SGD.
The gradient estimation in SubZero is applicable to parameter matrices, while LLMs mainly consist of dense layers. For other trainable parameters, such as biases and layer normalization parameters, we recommend using the gradient estimation in MeZO (Malladi et al., 2023), as these layers contain fewer parameters.
We introduce two useful strategies to implement our SubZero efficiently in memory.
In-place Operation. As indicated in Eqn. (7), directly computing the loss difference $\rho$ requires twice the memory of inference, as it must store both the parameter matrix set $\mathcal{W}$ and the perturbation matrix set $\tilde{\mathcal{Z}}$. To mitigate this, we draw inspiration from MeZO and utilize in-place operations. By employing the random seed trick, we store a random seed to compute $\rho$ (see lines 9-12 in Algorithm 3 and Algorithm 2) and regenerate the low-dimensional perturbation matrices $\bm{Z}_{1},\bm{Z}_{2},\cdots,\bm{Z}_{l}$ (see line 15 in Algorithm 3). Consequently, the memory cost for fine-tuning with SubZero is nearly equivalent to that of inference (see Table 1 and Table 5).
Per-layer Weight Update. FO optimizers update all model parameters after BP by storing the entire gradients in memory. In contrast, ZO optimizers like SubZero calculate gradient estimates by first determining the loss value difference from two forward passes, then calculating the gradient estimate for each layer using this difference along with the layer’s perturbation. To reduce memory usage during training, we can implement the parameter update with optimizer.step() after calculating the gradient estimate for each layer.
SubZero significantly reduces GPU memory consumption with the two implementation strategies. It should note that we use the per-layer weight update strategy for MeZO in all experiments.
To simplify hyperparameter tuning, we employ a norm alignment trick, allowing SubZero to directly utilize hyperparameter settings, such as the learning rate, from MeZO (Malladi et al., 2023). For a random perturbation matrix ${\bm{Z}}\in\mathbb{R}^{m\times n}$, and its low-rank approximation is $\hat{{\bm{Z}}}={\bm{U}}{\bm{Z}}^{\prime}{\bm{V}}^{\mathsf{T}}$, where ${\bm{U}}\in\mathbb{R}^{m\times r}$, ${\bm{V}}\in\mathbb{R}^{n\times r}$, and ${\bm{Z}}^{\prime}\in\mathbb{R}^{r\times r}$. If ${\bm{Z}}$ and ${\bm{Z}}^{\prime}$ are Gaussian random matrices, and ${\bm{U}}$ and ${\bm{V}}$ are column-orthogonal matrices, then we have:
$\mathbb{E}[\left\|{\bm{Z}}\right\|_{F}]=\sqrt{\frac{m\times n}{r^{2}}}\mathbb{E}\left[\|\hat{{\bm{Z}}}\|_{F}\right].$
Define $\mu=\sqrt{\frac{m\times n}{r^{2}}}$. Let MeZO’s learning rate be $\eta$ and perturbation scale be $\varepsilon$. There are two equivalent approaches to obtain the perturbation for SubZero. The first approach involves multiplying the random low-dimensional perturbation matrix by $\mu$, with SubZero adopting MeZO’s hyperparameters directly: $\eta^{\prime}=\eta$ and $\varepsilon^{\prime}=\varepsilon$. The second approach keeps the random low-dimensional perturbation matrix fixed and sets SubZero’s learning rate and perturbation scale as follows:
$\eta^{\prime}=\eta\mu^{2},\varepsilon^{\prime}=\varepsilon\mu.$
We argue that norm alignment is crucial for SubZero, as changing the rank $r$ affects the norm of the gradient estimate, complicating the fine-tuning of the associated learning rate.
S-MeZO (Liu et al., 2024), a new ZO method, aims to improve MeZO’s performance and convergence speed. However, its source code and detailed layer-wise hyperparameter configurations have not been released. Yang et al. (2024) reproduce S-MeZO using a fixed sparsity ratio for each layer, selected based on the best overall result shown in Fig. 6 of their paper. So we perform S-MeZO with this non-official implementation code available at https://github.com/yifanycc/AdaZeta.
A.5 Proofs
In practice, SubZero employs smaller and layer-specific low-rank perturbation matrices instead of a large model-scale projection matrix. However, it is more convenient to prove SubZero’s properties using a model-scale projection. Fortunately, the following lemma shows that the low-rank perturbation matrix for each layer can be represented as a layer-scale projection matrix, which is column orthogonal. {lemma}[] Let $\tilde{{\bm{Z}}}={\bm{U}}{\bm{Z}}{\bm{V}}^{\mathsf{T}}$, where ${\bm{U}}\in\mathbb{R}^{m\times r},{\bm{Z}}\in\mathbb{R}^{r\times r},{\bm{V}}\in\mathbb{R}^{n\times r}$, and ${\bm{U}}^{\mathsf{T}}{\bm{U}}={\bm{V}}^{\mathsf{T}}{\bm{V}}={\bm{I}}_{r}$. Then we have ${\rm vec}(\tilde{{\bm{Z}}})={\bm{P}}{{\rm vec}}({\bm{Z}})$ and ${\bm{P}}^{\mathsf{T}}{\bm{P}}={\bm{I}}_{r^{2}}$, where ${\bm{P}}={\bm{V}}\otimes{\bm{U}}$.
Proof.
Since ${\rm vec}({\bm{U}}{\bm{Z}}{\bm{V}}^{\mathsf{T}})=({\bm{V}}\otimes{\bm{U}}){\rm vec}({\bm{Z}})$, we only need to show $({\bm{V}}\otimes{\bm{U}})^{\mathsf{T}}({\bm{V}}\otimes{\bm{U}})={\bm{I}}_{r^{2}}$. In fact
$({\bm{V}}\otimes{\bm{U}})^{\mathsf{T}}({\bm{V}}\otimes{\bm{U}})=({\bm{V}}^{\mathsf{T}}\otimes{\bm{U}}^{\mathsf{T}})({\bm{V}}\otimes{\bm{U}})=({\bm{V}}^{\mathsf{T}}{\bm{V}})\otimes({\bm{U}}^{\mathsf{T}}{\bm{U}})={\bm{I}}_{r}\otimes{\bm{I}}_{r}={\bm{I}}_{r^{2}}.$
The proof is completed. ∎
We can also demonstrate that the low-rank perturbation matrices across all layers can be represented as a model-scale projection matrix. We first give the following lemma. {lemma}[] Let a block diagonal matrix ${\bm{P}}={\rm bdiag}({\bm{P}}_{1},{\bm{P}}_{2},\cdots,{\bm{P}}_{l})$ and $\tilde{{\bm{z}}}_{i}={\bm{P}}_{i}{\bm{z}}_{i}$, where ${\bm{P}}_{i}^{\mathsf{T}}{\bm{P}}_{i}={\bm{I}}_{r^{2}}$ and $i=1,2,\dots,l$. Then we have $\tilde{{\bm{z}}}={\bm{P}}{\bm{z}}$, where $\tilde{{\bm{z}}}=[\tilde{{\bm{z}}}_{1}^{\mathsf{T}},\dots,\tilde{{\bm{z}}}_{l}^{\mathsf{T}}]^{\mathsf{T}}$, ${\bm{z}}=[{\bm{z}}_{1}^{\mathsf{T}}$, $\dots,{\bm{z}}_{l}^{\mathsf{T}}]^{\mathsf{T}}$ and ${\bm{P}}^{\mathsf{T}}{\bm{P}}={\bm{I}}_{lr^{2}}$.
Proof.
It is easy to check that $\tilde{{\bm{z}}}={\bm{P}}{\bm{z}}$. Besides, we have
${\bm{P}}^{\mathsf{T}}{\bm{P}}={\rm bdiag}({\bm{P}}_{1}^{\mathsf{T}},\dots,{\bm{P}}_{l}^{\mathsf{T}}){\rm bdiag}({\bm{P}}_{1},\dots,{\bm{P}}_{l})={\rm bdiag}({\bm{P}}_{1}^{\mathsf{T}}{\bm{P}}_{1},\dots,{\bm{P}}_{l}^{\mathsf{T}}{\bm{P}}_{l})={\bm{I}}_{lr^{2}}.$
The proof is completed. ∎
We may define ${\bm{P}}={\rm bdiag}(\bm{V}_{1}\otimes\bm{U}_{1},\bm{V}_{2}\otimes\bm{U}_{2},\cdots,\bm{V}_{l}\otimes\bm{U}_{l})$ that satisfies ${\bm{P}}^{\mathsf{T}}{\bm{P}}={\bm{I}}$, ${\bm{z}}=[{\rm vec}({\bm{Z}}_{1})^{\mathsf{T}},{\rm vec}({\bm{Z}}_{2})^{\mathsf{T}},\dots,{\rm vec}({\bm{Z}}_{l})^{\mathsf{T}}]^{\mathsf{T}}$, and $\tilde{{\bm{z}}}=[{\rm vec}(\tilde{{\bm{Z}}}_{1})^{\mathsf{T}},{\rm vec}(\tilde{{\bm{Z}}}_{2})^{\mathsf{T}},\dots,{\rm vec}(\tilde{{\bm{Z}}}_{l})^{\mathsf{T}}]^{\mathsf{T}}$. Then according to Lemma A.5, the perturbation vector of SubZero is $\tilde{\bm{z}}={\bm{P}}\bm{z}$, which is similar as existing random subspace methods in Eqn. (4), but with SubZero’s projection matrix being block diagonal and column orthogonal.
To prove Theorem 5 and Theorem 5, we first introduce some definitions and lemmas about Gaussian distribution.
Defination 1.
[…]
Defination 2.
Let $\bm{z}\sim\mathcal{N}(\bm{0},{\bm{I}}_{n})$. We say $x$ is a chi-square random variable with degrees of freedom $n$ (denote by $x\sim\chi^{2}(n)$), if $x=\|\bm{z}\|^{2}$.
{lemma}
[] Let $\bm{z}\sim\mathcal{N}(\bm{0},{\bm{I}}_{n})$. For any orthogonal $(n\times n)$-matrix ${\bm{Q}}$ and continuous function $f$, we have $\mathbb{E}_{\bm{z}}[f(\bm{z})]=\mathbb{E}_{\bm{z}}[f({\bm{Q}}\bm{z})]$.
{lemma}
[] If $x\sim\chi^{2}(n)$, then we have
$\mathbb{E}_{x}[x]=n,\quad\mathrm{Var}_{x}[x]=2n.$
{lemma}
[] (Nesterov & Spokoiny, 2017) Let $f\in C^{2,2}_{L_{2}}(\mathbb{R}^{n})$. Then for all ${\bm{x}},{\bm{y}}\in\mathbb{R}^{n}$, we have
$|f({\bm{y}})-f({\bm{x}})-\langle\nabla f({\bm{x}}),{\bm{y}}-{\bm{x}}\rangle-\frac{1}{2}\langle\nabla^{2}f({\bm{x}})({\bm{y}}-{\bm{x}}),{\bm{y}}-{\bm{x}}\rangle|\leq\frac{L_{2}}{6}\|{\bm{y}}-{\bm{x}}\|^{3}.$
{lemma}
[] (Nesterov & Spokoiny, 2017) Let $\bm{z}\sim\mathcal{N}(\bm{0},{\bm{I}}_{n})$. For $0\leq t\leq 2$, we have
$\mathbb{E}_{\bm{z}}[\|\bm{z}\|^{t}]\leq n^{t/2}.$
For $t\geq 2$, we have
$n^{t/2}\leq\mathbb{E}_{\bm{z}}[\|\bm{z}\|^{t}]\leq(n+t)^{t/2}.$
{lemma}
[] Let $\bm{z}\sim\mathcal{N}(\bm{0},{\bm{I}}_{n})$. For all ${\bm{y}}\in\mathbb{R}^{n}$, we have
$\mathbb{E}_{\bm{z}}[\|\langle{\bm{y}},\bm{z}\rangle\bm{z}\|^{2}]=(n+2)\|{\bm{y}}\|^{2}.$
Proof.
Note that for any orthogonal $(n\times n)$-matrix ${\bm{Q}}$, we have
$\|\langle{\bm{y}},{\bm{Q}}\bm{z}\rangle{\bm{Q}}\bm{z}\|^{2}=\|\langle{\bm{Q}}^{\mathsf{T}}{\bm{y}},\bm{z}\rangle\bm{z}\|^{2},\quad\|{\bm{Q}}^{\mathsf{T}}{\bm{y}}\|=\|{\bm{y}}\|.$
In accordance with Lemma 2, we can set ${\bm{y}}=[1,0,\dots,0]^{\mathsf{T}}$, and only need to prove $\mathbb{E}_{\bm{z}}[\|\langle{\bm{y}},\bm{z}\rangle\bm{z}\|^{2}]=n+2$. Equipped with Lemma 2, we get
$\mathbb{E}_{\bm{z}}[\|\langle{\bm{y}},\bm{z}\rangle\bm{z}\|^{2}]=\mathbb{E}_{\bm{z}}\left[\sum_{i=1}^{n}\bm{z}_{1}^{2}\bm{z}_{i}^{2}\right]=\sum_{i=1}^{n}\mathbb{E}_{\bm{z}}[\bm{z}_{1}^{2}\bm{z}_{i}^{2}]=\mathbb{E}_{\bm{z}_{1}}[\bm{z}_{1}^{4}]+\mathbb{E}_{\bm{z}_{1}}[\bm{z}_{1}^{2}]\sum_{i=2}^{n}\mathbb{E}_{\bm{z}}[\bm{z}_{i}^{2}]=n+2.$
The proof is completed. ∎
See 5
Proof.
a) Evidently, the conclusion is established based on Lemma A.5 and Lemma A.5.
b)
Let $a_{{\bm{z}}}(\tau)=f({\bm{x}}+\tau{\bm{z}})-f({\bm{x}})-\tau\langle\nabla f({\bm{x}}),{\bm{z}}\rangle-\frac{\tau^{2}}{2}\langle\nabla^{2}f({\bm{x}}){\bm{z}},{\bm{z}}\rangle$. Lemma 2 implies that
$|a_{{\bm{z}}}(\pm\varepsilon)|\leq\frac{\varepsilon^{3}}{6}L_{2}\|{\bm{z}}\|^{3}.$
Note that
$\mathbb{E}_{\bm{z}}[\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})]-{\bm{P}}{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})$
[…]
Therefore, in accordance with Lemma 2, we have
$\|\mathbb{E}_{\bm{z}}[\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})]-{\bm{P}}{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|$
[…]
The proof is completed. ∎
See 5
Proof.
It is easy to check that $\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})={\bm{P}}\langle{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}}),\bm{z}\rangle\bm{z}$. Thus we have $\mathbb{E}_{\bm{z}}[\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})]={\bm{P}}{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})$. Combined with Lemma 2, we get $\mathbb{E}_{\bm{z}}[\|\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})\|^{2}]=(q+2)\|{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|^{2}$. Note that for any orthogonal $(q\times q)$-matrix ${\bm{Q}}$, we have
$\mathbb{E}_{\bm{z}}\left[\frac{\langle\nabla f({\bm{x}}),\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})\rangle^{2}}{\|{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|^{2}\|\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})\|^{2}}\right]$ $=\mathbb{E}_{\bm{z}}\left[\frac{\langle{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}}),\bm{z}\rangle^{2}}{\|{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|^{2}\|\bm{z}\|^{2}}\right]$
$=\mathbb{E}_{\bm{z}}\left[\frac{\langle{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}}),{\bm{Q}}\bm{z}\rangle^{2}}{\|{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|^{2}\|{\bm{Q}}\bm{z}\|^{2}}\right]$
$=\mathbb{E}_{\bm{z}}\left[\frac{\langle{\bm{Q}}^{\mathsf{T}}{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}}),\bm{z}\rangle^{2}}{\|{\bm{Q}}^{\mathsf{T}}{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|^{2}\|\bm{z}\|^{2}}\right].$
In accordance with Lemma 2, we can set ${\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})=[1,0,\dots,0]^{\mathsf{T}}$. Thus we have
$\mathbb{E}_{\bm{z}}\left[\frac{\langle\nabla f({\bm{x}}),\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})\rangle^{2}}{\|{\bm{P}}^{\mathsf{T}}\nabla f({\bm{x}})\|^{2}\|\hat{g}_{\varepsilon}({\bm{x}},{\bm{P}},\bm{z})\|^{2}}\right]=\mathbb{E}_{\bm{z}}\left[\frac{\bm{z}_{1}^{2}}{\|\bm{z}\|^{2}}\right]=\frac{1}{q}.$
The proof is completed. ∎
{lemma}
[] Let $h({\bm{y}})=f({\bm{x}}+{\bm{P}}{\bm{y}})$, where $f\in C^{1,1}_{L_{1}}(\mathbb{R}^{d})$ and $f$ is convex, and ${\bm{P}}^{\mathsf{T}}{\bm{P}}={\bm{I}}$, then we have $h\in C^{1,1}_{L_{1}}(\mathbb{R}^{q})$ and $h$ is convex.
Proof.
Note that convexity is an affine-invariant property (Nesterov, 2018), if $f$ is convex, we can obtain that $h$ is also convex.
The following proves that if $f$ is first $L_{1}$-smooth, then $h$ is also first $L_{1}$-smooth. For any ${\bm{y}}_{1}\in\mathbb{R}^{q}$ and ${\bm{y}}_{2}\in\mathbb{R}^{q}$, we have
$\left\|\nabla h({\bm{y}}_{1})-\nabla h({\bm{y}}_{2})\right\|$ $=\left\|{\bm{P}}^{\mathsf{T}}\nabla(f({\bm{x}}+{\bm{P}}{\bm{y}}_{1})-\bm{P}^{\mathsf{T}}\nabla(f({\bm{x}}+{\bm{P}}{\bm{y}}_{2})\right\|$
$\leq\left\|{\bm{P}}^{\mathsf{T}}\right\|\left\|\nabla(f({\bm{x}}+{\bm{P}}{\bm{y}}_{1})-\nabla(f({\bm{x}}+{\bm{P}}{\bm{y}}_{2})\right\|$
$\leq L_{1}\left\|{\bm{P}}({\bm{y}}_{1}-{\bm{y}}_{2})\right\|$
$=L_{1}\left\|{\bm{y}}_{1}-{\bm{y}}_{2}\right\|$
The proof is completed.
See 5
Proof.
In accordance with Lemma 2, we can transform the original problem $f\in C^{1,1}_{L_{1}}(\mathbb{R}^{d})$ into $h\in C^{1,1}_{L_{1}}(\mathbb{R}^{q})$ through affine transformation $h({\bm{y}})=f({\bm{x}}+{\bm{P}}{\bm{y}})$. The subsequent convergence proof can directly refer to Theorem 8 in (Nesterov & Spokoiny, 2017).
The proof is completed. ∎

### v3

@@ 6.1 Results under Different Experimental Settings
We also present several training loss curves in Fig. 2 (a)-(c), demonstrating that SubZero generally converges faster and achieves lower losses compared to MeZO.
6.2 Results on Non-Differentiable Objectives
Following MeZO [40], we apply SubZero to fine-tune RoBERTa-large and OPT-13B using full-parameter fine-tuning with two non-differentiable objectives—optimizing accuracy in classification tasks and F1 in the SQuAD task. As baselines, we also report results using the differentiable cross-entropy objective with Adam, MeZO, and SubZero. As shown in Table 6, SubZero consistently outperforms MeZO with the two non-differentiable objectives. The training loss curves in Fig. 2 (d) also demonstrate the advantage of SubZero over MeZO.
@@ 6.3 Memory Usage and Wall-Clock Time Analysis
We compare the memory consumption and wall-clock time of ZO methods (MeZO and SubZero), SGD, and inference-only approaches (zero-shot and in-context learning (ICL)) using OPT-13B (see Table 10 in Appendix 8.1). Since inference-only methods do not involve fine-tuning, they have zero wall-clock time and their memory usage reflects only the inference load. For fine-tuning, all methods were run for 20K steps. The ZO methods, including SubZero, achieved over a 1.8× reduction in memory usage compared to SGD. Notably, SubZero’s memory footprint closely aligns with MeZO’s, while offering improved performance. We use per-layer weight updates for MeZO and SubZero (see Appendix 8.2), resulting in nearly identical memory usage for FT and LoRA schemes when one decimal place is reserved.
Although SubZero introduces additional computational overhead for generating projection matrices via QR decomposition, the observed extra time overhead remains below 9% across all OPT model sizes, with a mean value of 4.79% of the total wall-clock time (see Table 11 in Appendix 8.1).
Note that due to differences in how steps are defined between ZO methods and SGD, direct wall-clock time comparisons between the two are not entirely meaningful.
@@ 6.4 Ablation Study
We conducted a thorough investigation of the effectiveness of our techniques. Table 5 shows that using a column-orthogonal projection matrix significantly outperforms a Gaussian random projection matrix, primarily due to the low-rank structure of the perturbation matrices (see experimental settings in Appendix 8.4). This low-rank perturbation is key to improving the quality of gradient estimation.
Next, Table 8 explores the effects of subspace rank $r$ and update frequency $F$ in Algorithm 3. The results demonstrate that SubZero is robust to variations in the subspace rank. However, performance drops sharply when the update frequency is too low, as the optimization becomes constrained to a single subspace for too long, limiting its adaptability.
Table 7: Subspace change frequency $F$ and rank $r$.
$F$$r$ 32 64 128
@@
Due to limited space, the ablation studies of SubZero on random seed, batch size, and combination with Adam are given in Appendix 8.1.
@@ 7 Conclusion
We have demonstrated that SubZero effectively fine-tunes large LLMs across various tasks and schemes with a memory cost comparable to that of inference. Extra experiments indicate that SubZero can optimize non-differentiable objectives. Our theory explains how SubZero reduces the variance of gradient estimates and hence accelerates convergence.
@@
Limitation. In addition to the representative first-order and primitive zero-order optimizers, we have yet to investigate the combinations of SubZero with other first-order and zero-order optimizers to evaluate the implications on convergence speed. While SubZero is compatible with memory-efficient techniques like parameter quantization [35], we have not thoroughly explored the practical effects of these combinations. A theoretical analysis of the reshaping strategy is valuable but left for future work.
Acknowledgments
This work is supported by the NSF of China (grant nos. 62131003 and 62102034), Beijing Municipal Science and Technology Project (Z241100001324011), the Ministry of Education, Singapore, under its AcRF Tier 2 Funding (Proposal ID: T2EP20224-0048). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.